OpenStudy (anonymous):

can anyone help? Use the quadratic formula to solve 2-5x=-3x^2. which values can be used for the values of a,b,and c

OpenStudy (anonymous):

2 - 5x = -3x^2 +3x^2 +3x^2 3x^2 -5x + 2 = 0 ax^2+bx+c a= 3 b= -5 c= 2

OpenStudy (owlcoffee):

In order o use the quadratic formula, you'll have to accomodate the equation into the form: \[ax^2+bx+c=0 \] Which in this case mans, leave a zero on either side of the equal sign, so: \[2-5x=-3x^2 \iff 3x^2-5x+2=0 \] with the form now accomodated, the solution can be found using the general equation.

OpenStudy (anonymous):

thank you thank you :)

OpenStudy (anonymous):

anytime ;)

OpenStudy (anonymous):

2 more? solve x^2-2x+5=0 by using the quadratic formula. what is the solution set

OpenStudy (anonymous):

the discriminant of a quadratic equation is -5. which answer describes the number and type of solutions of the equation? 2 complex solutions 2 real solutions 1 real solution 1 complex solution

OpenStudy (anonymous):

@rainbowswirl